I do not know if this is the right section or not, but please help me solve this equation: 6r/12 = -1/14
Answer by checkley75(3666)

6r/12 = -1/14
6r*14=12*-1
84r=-12
r=-12/84
r=-1/7 ans.
Proof:
(6*-1/7)/12=-1/14
(-6/7)/12=-1/14
-6/7*1/12=-1/14
-6/84=-1/14
-1/14=-1/14